1. Juan takes 6.3 hours with an average vleocity of 96 km/hr west to drive from El Paso, Tx. to Tucson, Az. What is Juan's displacement?

2. A truck traveling at 9.0 m/s north accelerates at the rate of +0.65 m/s2 for a total of 5.1 seconds. What is the velocity of the truck at the end of this time interval?

3. A boat is traveling in a straight line with a velocity of 6.0 m/s east. After an acceleration of 0.59 m/s2, the boat's velocity is 9.5 m/s. In what time interval did the acceleration occur?
1
Expert's answer
2020-12-06T17:26:51-0500

1)


"s=vt=(96)(6.3)=605\ km\text{ West}"

2)


"v_f=v_i+at=9+0.65(5.1)=12.3\frac{m}{s}"

3)


"t=\frac{v_f-v_i}{a}=\frac{9.5-6}{0.59}=5.9\ s"